# Diffusion Generative Recommendations with Continuous Tokens

### v3

@@ 4.1.4. Configurations
We implement all models using Python 3.11 and PyTorch 2.5.1 on an NVIDIA A800 (80 GB) GPU. We use AdamW with learning rates and weight decays of 0.00001/0.0001 for the backbone and 0.0001/0.001 for the tokenizer. The maximum length of the item sequence is configured to 20. The masking ratio $\rho$ is searched from 0 to 1 in 0.2 increments, while the number of sub-encoders is selected from {1, 2, 3, 4, 5}. Taking into account both performance and computational resources, we opt to utilize the Llama-3.2-1B-Instruct [^4] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct as our LLM backbone, which is renowned as one of the most popular pre-trained language models globally. The LLM backbone operates with a default temperature of 0. We create ten prompt templates and three response templates for the input and output of the LLM backbone, respectively. A random selection from these templates is made in each batch to enhance the model’s capacity to generalize across various template structures. The two loss weights $\gamma_{1}$ and $\gamma_{2}$ are sequential searched in the range of {0, 0.5, 1.0, 1.5, 2.0}. The basic embeddings for users and items are derived from both textual information encoded by SentenceBERT (Reimers and Gurevych, 2019) and collaborative knowledge learned through LightGCN (He et al., 2020). Our noise schedule consists of 1000 steps during training; during inference, it is resampled with fewer steps (100). We fix the unconditional training probability $\zeta$ as 0.1 suggested by (Ho and Salimans, [n. d.]), and the diffusion temperature $\tau$ to 0.5 following (Wang and He, 2025). We search for the personalized guidance strength $\omega$ in the range of {0, 2, 4, 6, 8, 10}. More implementation details are presented in Appendix C.
@@ Appendix A Related Work
The rapid advancement of Large Language Models (LLMs), such as ChatGPT and GPT-4, has significantly transformed natural language processing (Ning et al., 2025; Fan et al., 2025b; Qu et al., 2024a). These models, with their billion-scale parameters, excel in language comprehension, generation, and generalization, allowing them to adapt to new tasks and domains (Kaplan et al., 2020). As a result, LLMs have increasingly been integrated into recommender systems, enhancing their capabilities (Wang et al., 2025a; Zhao et al., 2024; Huang et al., 2025). To effectively leverage LLMs in recommendation tasks, two primary challenges must be addressed: the tokenization of users and items, and the generation of recommendations. Most research has focused on adapting recommender systems to the language domain using discretization techniques, which align with the discrete nature of natural language. A notable example is P5 (Geng et al., 2022), which unifies various recommendation tasks into natural language processing through multi-task prompt-based pre-training, achieving impressive zero-shot generalization with personalized prompts. Following this, simple tokenization methods like title indexing, independent indexing, and sequential indexing have become common (Hua et al., 2023), as seen in studies like TallRec (Bao et al., 2023) and LLMRec (Wei et al., 2024). However, these methods struggle to capture complex collaborative knowledge and are limited by LLMs’ input length constraints. To overcome these limitations, researchers have developed robust tokenizers that convert users, items, and their interactions into discrete tokens. Examples include TIGER (Rajput et al., 2023) and TokenRec (Qu et al., 2025), which use vector-quantized tokenizers to encode textual information and collaborative representations from graph neural network-based methods (He et al., 2020). Despite the performance improvements offered by advanced discrete-valued tokenizers, they are hindered by information loss during compression or quantization. Consequently, there is growing interest in continuous-valued tokens, which have shown promise in representation learning (Yue et al., 2023). However, this approach is still in its early stages of development. For instance, models like CoLLM (Zhang et al., 2025) and LlaRA (Liao et al., 2024) use continuous item representations to index items in LLM-based recommender systems through a projection layer that aligns collaborative and LLM token embeddings. However, such a direct approach faces challenges in bridging the substantial gap between the continuous collaborative knowledge that GNNs capture from user-item interactions and the semantic information that LLMs learn from discrete natural language.
A.2. Diffusion Models with Continuous Tokens.
@@
Denoising diffusion probabilistic models, also known as DDPM or diffusion models, function based on the concept of learning to reverse a gradual noise-addition process (Ho et al., 2020; Nichol and Dhariwal, 2021). In contrast to language models that commonly operate with discrete tokens, diffusion models work with continuous vectors, rendering them well-suited for tasks that involve continuous data, such as images. For example, palette (Saharia et al., 2022) demonstrates strong performance across a range of image-to-image generation tasks by employing a general diffusion model, eliminating the need for task-specific fine-tuning. Given the success of diffusion models in continuous data generation, an increasing number of studies have proposed combining large language models with diffusion models in cross-modal modeling, with both continuous and discrete data as inputs and targets. Examples include VAE-MAR (Li et al., 2024) and Next-Token Diffusion (Sun et al., 2024), which employ a variational autoencoder (VAE) to represent continuous image data as latent vectors and introduce denoising diffusion for autoregressive generation of these vectors. In these studies, language models continue their imposition on discrete language generation and comprehension, while diffusion models are incorporated as generation heads to operate within continuous domains. Despite that, the integration of diffusion models with auto-regressive models is currently in its nascent phase.
@@ Appendix B Further Explanation of ContRec
B.1. Prompt Templates
Inspired by the success of in-context learning in large language models, we employ two widely used prompting techniques, namely role play and few-shot learning. Let $\left\langle\cdot\right\rangle$” represents a continuous encoded token for users or items. A detailed example of our prompt template is outlined in the colorbox. Please refer to our code for the detailed prompts.
[…]
B.2. Algorithm of the Proposed Framework
The pseudo-code of the proposed framework during training and inference is described in Algorithm 1. The ultimate goal is to develop a comprehensive understanding of user preferences for generating effective and accurate recommendations. Top-K relevant items can be retrieved from the ranking list as the recommended output.
@@ B.2. Algorithm of the Proposed Framework
B.3. Derivation of Dispersive Loss
In our case, we derive the dispersive loss (Wang and He, 2025) from the typical contrastive learning loss, namely InfoNCE loss. Formally, the original InfoNCE loss (Oord et al., 2018) can be interpreted as a categorical cross-entropy objective that encourages high similarity between positive pairs and low similarity between negative pairs, as shown in Eq. (18). We can rewrite this equation as:
@@ Appendix C Further Analysis of Experimental Results
C.1. Discussion on Inference Costs.
As shown in Table 4, the majority of our model’s inference latency is attributed to the LLM’s reasoning phase, which accounts for approximately 88.60% of the total inference time across all four datasets. Nevertheless, in line with the “deep thinking” paradigm adopted by recent advanced large language models such as DeepSeek-R1 and GPT-5, we suggest that this ”latency” is essential, as it provides transparency into how the model generates recommendations and thereby fosters user trust. Given these considerations, we propose that the most suitable application scenarios for our model are in personalized conversational recommendation systems with user interfaces, rather than as a back-end algorithm for large-scale online recommendation tasks involving billions of web requests.
Table 4. Average inference time (seconds) per batch.
Dataset LlaRA TIGER TokenRec DeftRec LLM Diffusion
LastFM 0.4255 0.3502 0.0240 1.6110 1.5244 0.0866
ML1M 0.7370 0.5990 0.0367 2.4264 2.0994 0.3271
Games 0.5281 0.4316 0.0273 1.7338 1.4920 0.2418
Beauty 0.5423 0.4388 0.0292 1.7643 1.5385 0.2258
C.2. Analysis of Generalizability.
To assess the Generalizability of our approach and its LLM-based counterparts, we utilize the Beauty and Games datasets. In these datasets, we exclude the 5% of users with the least interaction history from the training split, treating them as newly added users in the evaluation set. The results demonstrate that advanced LLM-based methods (i.e., TokenRec, TIGER, LLaRA, and ContRec) maintain strong performance in this scenario. This can be attributed to their ability to leverage textual information (e.g., title and category) as supplementary cues for inferring user preferences when interaction history is limited. Furthermore, our hybrid retrieval process encourages ContRec to predict items with relevant categories and brands, making it particularly effective in cases involving previously unseen users.
Model Beauty Games
HR@10 NDCG@10 HR@10 NDCG@10
P5 0.0255 0.0156 0.0643 0.0368
CoLLM 0.0263 0.0171 0.0663 0.0373
TokenRec 0.0374 0.0197 0.0747 0.0414
TIGER 0.0384 0.0195 0.0759 0.0429
LlaRA 0.0380 0.0198 0.0770 0.0436
ContRec 0.0391 0.0205 0.0787 0.0450
C.3. Implementation Details.
The proposed model is developed based on Hugging Face and PyTorch, leveraging multiple NVIDIA A800 GPUs (80GB). We employ the meta-Llama/Llama-3.2.-1B-Instruct model as the foundation of our LLM, which is subsequently quantized into a 4-bit format to facilitate efficient fine-tuning. In ContRec, the trainable parameters are divided into three parts, with their respective sizes detailed in Table 6. Additionally, the total model parameter size is estimated to be 3,058.967 MB. To maximize the utilization of our computing hardware, the NVIDIA A800 GPU (80GB), we have configured a batch size of 24 and a maximum of 200 epochs. The training process on the four datasets (ML1M, Games, Beauty, and LastFM) demands considerable time, exceeding 72 hours, 60 hours, 50 hours, and 24 hours on a single A800 GPU, respectively. For all baselines, we conduct negative sampling from the uniform distribution at the ratio of 1:1, which is not conducted in DreamRec and ContRec.
Table 6. Statistics of the trainable parameters.
Name Type Params
1 model PeftModelForCausalLM 760 M
2 diffusion_net Transformer 5.8 M
3 item_tokenizer MLP 2.6 M
4 user_tokenizer MLP 2.6 M
5 condition_net MLP 1.1 M
C.4. Analysis of Success and Failure Cases.
Table 7 shows two success cases (users 2715 and 36183) and a failure case (user 7849) of ContRec. Notably, the failure case for user 7849 demonstrates a scenario in which ContRec makes an incorrect prediction due to the user’s shift in type, platform, and category preferences. This highlights a potential limitation of our model: the LLM-based reasoning tends to favor items with categories similar to those in the user’s historical interactions, making it challenging to accommodate sudden shifts in user preferences.

### v4

@@ 4.1.4. Configurations
We implement all models using Python 3.11 and PyTorch 2.5.1 on an NVIDIA A800 (80 GB) GPU. We use AdamW with learning rates and weight decays of 0.00001/0.0001 for the backbone and 0.0001/0.001 for the tokenizer. The maximum length of the item sequence is configured to 20. The masking ratio $\rho$ is searched from 0 to 1 in 0.2 increments, while the number of sub-encoders is selected from {1, 2, 3, 4, 5}. Taking into account both performance and computational resources, we opt to utilize the Llama-3.2-1B-Instruct [^4] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct as our LLM backbone, which is renowned as one of the most popular pre-trained language models globally. The LLM backbone operates with a default temperature of 0. We create ten prompt templates and three response templates for the input and output of the LLM backbone, respectively. A random selection from these templates is made in each batch to enhance the model’s capacity to generalize across various template structures. The two loss weights $\gamma_{1}$ and $\gamma_{2}$ are sequential searched in the range of {0, 0.5, 1.0, 1.5, 2.0}. The basic embeddings for users and items are derived from both textual information encoded by SentenceBERT (Reimers and Gurevych, 2019) and collaborative knowledge learned through LightGCN (He et al., 2020). Our noise schedule consists of 1000 steps during training; during inference, it is resampled with fewer steps (100). We fix the unconditional training probability $\zeta$ as 0.1 suggested by (Ho and Salimans, [n. d.]), and the diffusion temperature $\iota$ to 0.5 following (Wang and He, 2025). We search for the personalized guidance strength $\omega$ in the range of {1, 2, 4, 6, 8, 10}. More implementation details are presented in Appendix C.
@@ 5. Conclusion
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (project No.62406336). The research described in this paper has been partially supported by the General Research Funds from the Hong Kong Research Grants Council (project no. PolyU 15207322, 15200023, 15206024, and 15224524), internal research funds from Hong Kong Polytechnic University (project no. P0042693, P0048625, and P0051361), and Sheertek International (HK) Limited. This work was supported by computational resources provided by The Centre for Large AI Models (CLAIM) of The Hong Kong Polytechnic University.
@@ Appendix A Related Work
LLM-based Recommender Systems. The rapid advancement of Large Language Models (LLMs), such as ChatGPT and GPT-4, has significantly transformed natural language processing (Ning et al., 2025; Fan et al., 2025b; Qu et al., 2024; Kaplan et al., 2020). As a result, LLMs have increasingly been integrated into recommender systems, enhancing their capabilities (Wang et al., 2025a; Zhao et al., 2024; Huang et al., 2025). To effectively leverage LLMs in recommendation tasks, two primary challenges must be addressed: the tokenization of users and items, and the generation of recommendations. Most research has focused on adapting recommender systems to the language domain using discretization techniques, which align with the discrete nature of natural language. A notable example is P5 (Geng et al., 2022), which unifies various recommendation tasks into natural language processing through multi-task prompt-based pre-training, achieving impressive zero-shot generalization with personalized prompts. Following this, simple tokenization methods like title indexing, independent indexing, and sequential indexing have become common (Hua et al., 2023), as seen in studies like TallRec (Bao et al., 2023) and LLMRec (Wei et al., 2024). However, these methods struggle to capture complex collaborative knowledge and are limited by LLMs’ input length constraints. To overcome these limitations, researchers have developed robust tokenizers that convert users, items, and their interactions into discrete tokens. Examples include TIGER (Rajput et al., 2023) and TokenRec (Qu et al., 2025), which use vector-quantized tokenizers to encode textual information and collaborative representations from graph neural network-based methods (He et al., 2020). Despite the performance improvements offered by advanced discrete-valued tokenizers, they are hindered by information loss during compression or quantization. Consequently, there is growing interest in continuous-valued tokens, which have shown promise in representation learning (Yue et al., 2023; Zhang et al., 2025a; Liao et al., 2024). However, this approach is still in its early stages of development. Such a direct approach faces challenges in bridging the substantial gap between the continuous collaborative knowledge that GNNs capture from user-item interactions and the semantic information that LLMs learn from discrete natural language.
@@
Diffusion Models with Continuous Tokens. Denoising diffusion probabilistic models, also known as DDPM or diffusion models, function based on the concept of learning to reverse a gradual noise-addition process (Ho et al., 2020; Nichol and Dhariwal, 2021). In contrast to language models that commonly operate with discrete tokens, diffusion models work with continuous vectors, rendering them well-suited for tasks that involve continuous data, such as images. For example, palette (Saharia et al., 2022) demonstrates strong performance across a range of image-to-image generation tasks by employing a general diffusion model, eliminating the need for task-specific fine-tuning. Given the success of diffusion models in continuous data generation, an increasing number of studies have proposed combining large language models with diffusion models in cross-modal modeling, with both continuous and discrete data as inputs and targets. Examples include VAE-MAR (Li et al., 2024) and Next-Token Diffusion (Sun et al., 2024), which employ a variational autoencoder (VAE) to represent continuous image data as latent vectors and introduce denoising diffusion for autoregressive generation of these vectors. In these studies, language models continue their imposition on discrete language generation and comprehension, while diffusion models are incorporated as generation heads to operate within continuous domains. Despite that, the integration of diffusion models with auto-regressive models is currently in its nascent phase.
@@ Appendix B Further Explanation of ContRec
Algorithm of the Proposed Framework. The pseudo-code of the proposed framework during training and inference is described in Algorithm 1. The ultimate goal is to develop a comprehensive understanding of user preferences for generating effective and accurate recommendations. Top-K relevant items can be retrieved from the ranking list as the recommended output.
@@
Derivation of Dispersive Loss. In our case, we derive the dispersive loss (Wang and He, 2025) from the typical contrastive learning loss, namely InfoNCE loss. Formally, the original InfoNCE loss (Oord et al., 2018) can be interpreted as a categorical cross-entropy objective that encourages high similarity between positive pairs and low similarity between negative pairs, as shown in Eq. (18). We can rewrite this equation as:
@@
Prompt Templates. Inspired by the success of in-context learning in large language models, we employ two widely used prompting techniques, namely role play and few-shot learning. Let $\left\langle\cdot\right\rangle$” represents a continuous encoded token for users or items. A detailed example of our prompt template is outlined in the colorbox. Please refer to our code for the detailed prompts.
[…]
@@ Appendix C Further Analysis of Experimental Results
Analysis of Success and Failure Cases. Table 4 shows two success cases (users 2715 and 36183) and a failure case (user 7849) of ContRec. Notably, the failure case for user 7849 demonstrates a scenario in which ContRec makes an incorrect prediction due to the user’s shift in type, platform, and category preferences. This highlights a potential limitation of our model: the LLM-based reasoning tends to favor items with categories similar to those in the user’s historical interactions, making it challenging to accommodate sudden shifts in user preferences.
@@
Discussion on Inference Costs. As shown in Table 5, the majority of our model’s inference latency is attributed to the LLM’s reasoning phase, which accounts for approximately 88.60% of the total inference time across all four datasets. Nevertheless, in line with the “deep thinking” paradigm adopted by recent advanced large language models such as DeepSeek-R1 and GPT-5, we suggest that this ”latency” is essential, as it provides transparency into how the model generates recommendations and thereby fosters user trust. Given these considerations, we propose that the most suitable application scenarios for our model are in personalized conversational recommendation systems with user interfaces, rather than as a back-end algorithm for large-scale online recommendation tasks involving billions of web requests.
Table 5. Average inference time (seconds) per batch.
Dataset LlaRA TIGER TokenRec ContRec LLM Diffusion
LastFM 0.4255 0.3502 0.0240 1.6110 1.5244 0.0866
ML1M 0.7370 0.5990 0.0367 2.4264 2.0994 0.3271
Games 0.5281 0.4316 0.0273 1.7338 1.4920 0.2418
Beauty 0.5423 0.4388 0.0292 1.7643 1.5385 0.2258
Analysis of Generalizability. To assess the Generalizability of our approach and its LLM-based counterparts, we utilize the Beauty and Games datasets. In these datasets, we exclude the 5% of users with the least interaction history from the training split, treating them as newly added users in the evaluation set. The results demonstrate that advanced LLM-based methods (i.e., TokenRec, TIGER, LLaRA, and ContRec) maintain strong performance in this scenario. This can be attributed to their ability to leverage textual information (e.g., title and category) as supplementary cues for inferring user preferences when interaction history is limited. Furthermore, our hybrid retrieval process encourages ContRec to predict items with relevant categories and brands, making it particularly effective in cases involving previously unseen users.
Model Beauty Games
HR@10 NDCG@10 HR@10 NDCG@10
P5 0.0255 0.0156 0.0643 0.0368
CoLLM 0.0263 0.0171 0.0663 0.0373
TokenRec 0.0374 0.0197 0.0747 0.0414
TIGER 0.0384 0.0195 0.0759 0.0429
LlaRA 0.0380 0.0198 0.0770 0.0436
ContRec 0.0391 0.0205 0.0787 0.0450
Implementation Details. The proposed model is developed based on Hugging Face and PyTorch, leveraging multiple NVIDIA A800 GPUs (80GB). We employ the meta-Llama/Llama-3.2.-1B-Instruct model as the foundation of our LLM, which is subsequently quantized into a 4-bit format to facilitate efficient fine-tuning. In ContRec, the trainable parameters are divided into three parts, with their respective sizes detailed in Table 7. Additionally, the total model parameter size is estimated to be 3,058.967 MB. To maximize the utilization of our computing hardware, the NVIDIA A800 GPU (80GB), we have configured a batch size of 24 and a maximum of 200 epochs. The training process on the four datasets (ML1M, Games, Beauty, and LastFM) demands considerable time, exceeding 72 hours, 60 hours, 50 hours, and 24 hours on a single A800 GPU, respectively. For all baselines, we conduct negative sampling from the uniform distribution at the ratio of 1:1, which is not conducted in DreamRec and ContRec.
Table 7. Statistics of the trainable parameters.
Name Type Params
1 model PeftModelForCausalLM 760 M
2 diffusion_net Transformer 5.8 M
3 item_tokenizer MLP 2.6 M
4 user_tokenizer MLP 2.6 M
5 condition_net MLP 1.1 M